OpenStudy (anonymous):

HELP ASAP. 3(c-2)=2(c-6) A. 2 B. -6 C. no solution D. identity

OpenStudy (tkhunny):

Have you considered the Distributive Property to eliminate the parentheses?

OpenStudy (jessiegonzales):

B.-6

OpenStudy (anonymous):

thank you!

OpenStudy (tkhunny):

3(c-2)=2(c-6) Here is what you SHOULD have done. 3c - 6 = 2c - 12 c - 6 = -12 c = -6 @jessiegonzales will not be there on your exam.

OpenStudy (jessiegonzales):

sorry.. didn't need to give me a warning

OpenStudy (tkhunny):

Yes, I did. That's exactly what that warning is for. Just don't do it. It was only a warning. Learn from it and move on.